If sin x is G.M of sin y and cos y, then find the value of cos 2x.
 
Ans.  2cos2(pi/4 + x)

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
sin²x = sin y cosy
 
Now
 
Co2 x = 1 - 2 sin²x = 1 - 2 siny sosy = 1 - sin2y
 
I think you have provided some incorrect information